# Space spanned by characteristic exponents

### v2

@@ 5.5. Complex conjugation and absolute value
By Lemma 6.2, there exist $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}(f_{c})(\overline{L})$ and $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{L}\setminus S$ such that $((\mathbf{n}_{L}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangle $(\mathbb{D}(\mathcal{O}_{L,S})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(L)_{\mathbb{Q}}$. For every $i\in\mathbb{Z}_{>0}$, we can take $y_{i}\in{\rm Per}(f)(\overline{K})$ such that $x_{i}\in(y_{i})_{c}$, and hence
@@
After removing finitely many terms, we can assume that $y_{i}\notin W_{c}\cup W_{\tau(c)}$ for all $i\geq 1$. It follows that $\mathbf{n}_{K}(\lambda(y_{i}))\in{\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ for $i\geq 1$. Observe that $(\phi_{v_{i}}\circ s_{p})_{i=1}^{\infty}$ is a $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-sequence. It follows that
@@
is an upper triangle $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(K)_{\mathbb{Q}}$. Since
@@
1. The case $f$ is PCF. According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. We conclude that $f$ is defined over $\overline{\mathbb{Q}}$ since it is non-exceptional, and Theorem 1.4 was already proved in §4.
2. The case $f$ is not PCF. Then Theorem 1.4 is a direct consequence of Theorem 5.7.
We finish the proof of Theorem 1.4.∎
@@ 7. Proofs of the Applications
Recall that our goal is to prove the Zariski-dense orbit conjecture for $f=f_{1}\times\cdots\times f_{N}:X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}\to X$ where $f_{1},\dots,f_{N}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ are non-constant endomorphisms on the projective line over an algebraically closed field $k$ of characteristic 00.
Without loss of generality, we may assume that $k$ is of finite transcendence degree over $\mathbb{Q}$. Fix an embedding of $k$ into $\mathbb{C}$. We view $f$ as an endomorphism on $X$ defined over $\mathbb{C}$. According to [29, Theorem 3.34], we may assume that all $f_{j}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ has degree at least $2$ for $1\leq j\leq N$.
Assume first that all $f_{j}$ are not exceptional, $1\leq j\leq N$. Corollary 1.7 implies that we can take $x_{j}\in{\rm Per}^{*}(f_{j})(\mathbb{C})$ for $1\leq j\leq N$ such that the multipliers $\rho_{f_{1}}(x_{1}),\cdots,\rho_{f_{N}}(x_{N})$ are multiplicatively independent in $\mathbb{C}$. After replacing $f$ by an iterate, we may assume that $f_{j}(x_{j})=x_{j}$ for $1\leq j\leq N$, and the multipliers $(\rho_{f_{j}}(x_{j})=f_{j}^{\prime}(x_{j}))_{j=1}^{N}$ are still multiplicatively independent. Denote $x=(x_{1},\dots,x_{N})\in X(k)$. Then $x$ is a fixed point of $f$ (smooth in the fixed locus of $f$) such that the eigenvalues of $df|_{x}$ are nonzero and multiplicatively independent. Then the conclusion follows from [1].
Assume that all $f_{j}$ are exceptional, $1\leq j\leq N$. This case is easy, and we just refer to the proof in the first several paragraphs of [29, §9.3].
We may assume that $0\leq s\leq N$ such that $f_{1},\cdots,f_{s}$ are not exceptional and $f_{s+1},\cdots,f_{N}$ are exceptional. Let $l(f)=\min\{s,N-s\}\geq 0$. Then we have done in the case $l(f)=0$. Then an induction on $l(f)\in\mathbb{Z}_{\geq 0}$ proves Theorem 1.12, as shown in the last several paragraphs of [29, §9.3].∎
@@
Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $\geq 2$. We aim to study the condition $(1)$ $f$ is PCF. Our goal is to prove that $(1)$ is equivalent to each of the conditions $(2)$ and $(3)$ in Theorem 1.14, which are determined by the multiplier spectrum and the length spectrum of $f$, respectively.
@@
$(3)^{\prime}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $\mathbf{n}_{\mathbb{Q}}({\rm rog}(\lvert\rho_{f}(x)\lvert))$ for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
Now we prove that the three conditions $(1),(2)^{\prime}$, and $(3)^{\prime}$ are equivalent.
@@
Suppose that $f$ is PCF. By [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in $\mathcal{M}_{d}$, except for the family of flexible Lattès maps. If $f$ is flexible Lattès, then according to [18, Lemma 5.6], we have $\rho_{f}(x)\in\mathbb{Z}$ for all $x\in{\rm Per}(f)(\mathbb{C})$. If $f$ is defined over $\overline{\mathbb{Q}}$, then clearly $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})={\rm Per}(f)(\mathbb{Q})$. Thus, we always have $\rho_{f}(x),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
Suppose that $(2)^{\prime}$ is false, then
@@
By [18, Corollary 3.9], $f$ cannot be a flexible Lattès map, hence $f$ is defined over $\overline{\mathbb{Q}}$, and over a number field $K$. We use the notation and ideas from §6.1 in the case where $\tau=\mathrm{id}$. Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all the archimedean places such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$. For every $v\in\mathcal{M}_{K}\setminus B$, the reduction $f_{v}$ is still PCF and its critical orbits come from those of $f$. Arguing similarly as in §6.1, it is easy to see that the set
@@
is co-finite in ${\rm Per}^{*}(f)(\mathbb{C})$, i.e., ${\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}$ is finite. It is well-known that ${\rm rank}(\mathcal{O}_{K,B}^{\times})=|B|-1<\infty$ (cf. [21, Theorem 3.12]). Note that
@@
$\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{K}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}\leq|{\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}|+|B|-1<\infty,$
which implies $\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}<\infty$, contradicting the assumption. Thus $(2)^{\prime}$ must hold.
$(3)^{\prime}$ follows from an argument similar to the above paragraph, corresponding to the case where $\tau$ is the complex conjugation of §6.1.
$(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$:
Let $\tau$ be the identity on $\mathbb{C}$ in the case $(2)^{\prime}$ and be the complex conjugation in the case $(3)^{\prime}$. We will show $(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$ simultaneously. Suppose that $f$ is not PCF. In particularly, $f$ is not flexible Lattès. By $(2)^{\prime}$ (or $(3)^{\prime}$) and Theorem 1.2, the map $f$ is defined over $\overline{\mathbb{Q}}$, hence over a number field $K$. We use the notation and ideas from §6.1 for $\tau$. Arguing similarly as in §6.1, after enlarging the number field $K$, we can take a finite set $B\subseteq\mathcal{M}_{K}$ of places containing all the archimedean places satisfying the following conditions:
  • $K/\mathbb{Q}$ is Galois and $\tau(K)=K$, hence $\tau|_{K}\in{\rm Gal}(K/\mathbb{Q})$;
@@
A small modification of the proof of Lemma 6.2 shows that there exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ satisfying the following conditions:
@@
For $i\geq 1$, let $p_{i}$ be the rational prime below $v_{i}$ and $\widetilde{B}\subseteq\mathcal{M}_{\mathbb{Q}}$ be the set of places of $\mathbb{Q}$ lying below the places in $B$. From (7.15) and (7.16), it is easy to see that the pair $((\mathbf{n}_{\mathbb{Q}}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{p_{i}})_{i=1}^{\infty})$ is an upper triangle $(\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$. Note that $\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ is $\tau$-invariant, where $\tau$ descends to $\mathbb{D}(\mathbb{Q})$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$. By Corollary 5.5 and Remark 5.6, this contradicts $(2)^{\prime}$ and $(3)^{\prime}$, respectively. Thus, $f$ must be PCF.
@@ Acknowledgement
The second-named author Junyi Xie would like to thank Thomas Gauthier, Gabriel Vigny, Charles Favre, and Serge Cantat for helpful discussions. The authors would like to thank the anonymous referee for many thoughtful comments, which help to improve the manuscript.
The first-named author would like to thank Beijing International Center for Mathematical Research in Peking University for the invitation. The second and third-named authors Junyi Xie and Geng-Rui Zhang are supported by NSFC Grant (No.12271007).

### v3

@@ 5.5. Complex conjugation and absolute value
By Lemma 6.2, there exist sequences $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}(f_{c})(\overline{L})$ and $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{L}\setminus S$ such that $((\mathbf{n}_{L}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangular $(\mathbb{D}(\mathcal{O}_{L,S})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(L)_{\mathbb{Q}}$. For every $i\in\mathbb{Z}_{>0}$, we can take $y_{i}\in{\rm Per}(f)(\overline{K})$ such that $x_{i}\in(y_{i})_{c}$, and hence
@@
After removing finitely many terms, we may assume that $y_{i}\notin W_{c}\cup W_{\tau(c)}$ for all $i\geq 1$. It follows that $\mathbf{n}_{K}(\lambda(y_{i}))\in{\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ for $i\geq 1$. Observe that $(\phi_{v_{i}}\circ s_{p})_{i=1}^{\infty}$ is a $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-sequence. It follows that
@@
is an upper triangular $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(K)_{\mathbb{Q}}$. Since
@@
1. The case where $f$ is PCF. According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ within the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. Since $f$ is non-exceptional, it is defined over $\overline{\mathbb{Q}}$, and Theorem 1.4 was already proved in §4.
2. The case where $f$ is not PCF. Then Theorem 1.4 is a direct consequence of Theorem 5.7.
This completes the proof of Theorem 1.4.∎
@@ 7. Proofs of the Applications
Recall that our goal is to prove the Zariski-dense orbit conjecture for $f=f_{1}\times\cdots\times f_{N}:X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}\to X$, where $f_{1},\dots,f_{N}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ are non-constant endomorphisms on the projective line over an algebraically closed field $k$ of characteristic 0.
Without loss of generality, we may assume that $k$ is of finite transcendence degree over $\mathbb{Q}$. Fix an embedding of $k$ into $\mathbb{C}$. We view $f$ as an endomorphism on $X$ defined over $\mathbb{C}$. According to [29, Theorem 3.34], we may assume that all $f_{j}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ have degree at least $2$ for $1\leq j\leq N$.
Assume first that all $f_{j}$ are not exceptional, $1\leq j\leq N$. Corollary 1.7 implies that we can take $x_{j}\in{\rm Per}^{*}(f_{j})(\mathbb{C})$ for $1\leq j\leq N$ such that the multipliers $\rho_{f_{1}}(x_{1}),\dots,\rho_{f_{N}}(x_{N})$ are multiplicatively independent in $\mathbb{C}$. After replacing $f$ by an iterate, we may assume that $f_{j}(x_{j})=x_{j}$ for $1\leq j\leq N$, and the multipliers $(\rho_{f_{j}}(x_{j})=f_{j}^{\prime}(x_{j}))_{j=1}^{N}$ remain multiplicatively independent. Let $x=(x_{1},\dots,x_{N})\in X(k)$. Then $x$ is a fixed point of $f$ (smooth in the fixed locus of $f$) such that the eigenvalues of $df|_{x}$ are nonzero and multiplicatively independent. The conclusion then follows from [1].
If all $f_{j}$ are exceptional, $1\leq j\leq N$, this case is relatively easy, and we refer to [29, Theorem 4.1, Theorem 1.14, and Lemma 3.30] (see also [29, §9.3]).
Now assume that $0\leq s\leq N$ such that $f_{1},\dots,f_{s}$ are not exceptional and $f_{s+1},\dots,f_{N}$ are exceptional. Let $l(f)=\min\{s,N-s\}\geq 0$. The case $l(f)=0$ has just been treated. An induction on $l(f)N\in\mathbb{Z}_{\geq 0}$ then proves Theorem 1.12, as shown in the last several paragraphs of [29, §9.3].∎
@@
Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $\geq 2$. We aim to study the condition: (1) $f$ is PCF. Our goal is to prove that (1) is equivalent to each of the conditions (2) and (3) in Theorem 1.14, which are determined by the multiplier spectrum and the length spectrum of $f$, respectively.
@@
$(3)^{\prime}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $)\mathbf{n}_{\mathbb{Q}}({\rm rog}(\lvert\rho_{f}(x)\lvert$)) for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
We now prove that the three conditions (1), $(2)^{\prime}$, and $(3)^{\prime}$ are equivalent.
@@
Suppose that $f$ is PCF. By [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in $\mathcal{M}_{d}$, except for the family of flexible Lattès maps. If $f$ is flexible Lattès, then according to [19, Lemma 5.6], we have $\rho_{f}(x)\in\mathbb{Z}$ for all $x\in{\rm Per}(f)(\mathbb{C})$. If $f$ is defined over $\overline{\mathbb{Q}}$, then clearly $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})={\rm Per}(f)(\mathbb{Q})$. Thus, in both cases, $\rho_{f}(x),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
Suppose that $(2)^{\prime}$ is false, i.e.,
@@
By [19, Corollary 3.9], $f$ cannot be a flexible Lattès map, so $f$ is defined over $\overline{\mathbb{Q}}$, and hence over some number field $K$. We use the notation and ideas from §6.1 with $\tau=\mathrm{Id}$. Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all archimedean places such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$. For every $v\in\mathcal{M}_{K}\setminus B$, the reduction $f_{v}$ is still PCF, and its critical orbits come from those of $f$. Arguing similarly as in §6.1, it is easy to see that the set
@@
is co-finite in ${\rm Per}^{*}(f)(\mathbb{C})$, i.e., ${\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}$ is finite. It is well-known that ${\rm rank}(\mathcal{O}_{K,B}^{\times})=(\#B)-1<\infty$ (cf. [21, Theorem 3.12]). Note that
@@
$\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{K}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}\leq\#({\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W})+(\#B)-1$
is finite, which implies $\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}<\infty$, contradicting the assumption. Thus $(2)^{\prime}$ must hold.
$(3)^{\prime}$ follows from a similar argument corresponding to the case where $\tau$ is complex conjugation in §6.1.
$(2)^{\prime}$ or $(3)^{\prime}\Rightarrow$ (1):
Let $\tau$ be the identity on $\mathbb{C}$ in the case $(2)^{\prime}$ and complex conjugation in the case $(3)^{\prime}$. We prove $(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$ simultaneously. Suppose $f$ is not PCF. In particular, $f$ is not a flexible Lattès map. By $(2)^{\prime}$ (or $(3)^{\prime}$) and Theorem 1.2, the map $f$ is defined over $\overline{\mathbb{Q}}$, hence over some number field $K$. We use the notation and ideas from §6.1 for $\tau$. Arguing similarly as in §6.1, after enlarging the number field $K$, we can choose a finite set $B\subseteq\mathcal{M}_{K}$ of places containing all archimedean places satisfying the following conditions:
  • $K/\mathbb{Q}$ is Galois and $\tau(K)=K$, so $\tau|_{K}\in{\rm Gal}(K/\mathbb{Q})$;
@@
A slight modification of the proof of Lemma 6.2 shows that there exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ satisfying:
@@
For $i\geq 1$, let $p_{i}$ be the rational prime below $v_{i}$ and $\widetilde{B}\subseteq\mathcal{M}_{\mathbb{Q}}$ the set of places of $\mathbb{Q}$ lying below the places in $B$. From (7.15) and (7.16), it is easy to see that the pair $((\mathbf{n}_{\mathbb{Q}}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{p_{i}})_{i=1}^{\infty})$ is an upper triangular $(\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$. Note that $\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ is $\tau$-invariant, where $\tau$ descends to $\mathbb{D}(\mathbb{Q})$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$. By Corollary 5.5 and Remark 5.6, this contradicts $(2)^{\prime}$ and $(3)^{\prime}$, respectively. Thus, $f$ must be PCF. ∎
@@ Acknowledgement
The second-named author Junyi Xie would like to thank Thomas Gauthier, Gabriel Vigny, Charles Favre, and Serge Cantat for helpful discussions. The authors would like to thank the anonymous referee for many thoughtful comments, which help to improve the manuscript. The first-named author would like to thank Beijing International Center for Mathematical Research in Peking University for the invitation. The first-named author, Zhuchao Ji, is supported by National Key R&D Program of China (No.2025YFA1018300), NSFC Grant (No.12401106), and ZPNSF grant (No.XHD24A0201). The second- and third-named authors, Junyi Xie and Geng-Rui Zhang, are supported by NSFC Grant (No.12271007).